Prove that there does not exist a linear map T : R5 ! R5

such that range T = null T.


1
Expert's answer
2021-07-12T16:21:01-0400

By the rank-nullity theorem we have "\dim\operatorname{Im}T+\dim\ker T=\dim\mathbb R^5=5"

If "\operatorname{Im}T=\ker T", then "\dim\operatorname{Im}T=\dim\ker T=\frac{5}{2}". It is impossible, so "\operatorname{Im}T\neq\ker T"